Montserrat Volcano Observatory

Daily Report
Report for the period 6pm 11 May
to 6pm 12 May 1998

The level of volcanic activity has remained very low.

Between 4pm yesterday and 4pm today there were 20 volcano-tectonic earthquakes and 2 rockfall events recorded by the seismograph network. These latest VT's are located about 2-3km directly below the dome.

Dust levels were low today. Large amounts of sulphur dioxide were observed extending over the upper parts of the volcano. Work was begun to build for a new seismograph station at Brodericks and GPS surveying was continued on the island-wide network.

Residents are reminded that the volcano remains in a potentially active state, and that they should continue listening to Radio Montserrat in case the present quiet situation at the volcano changes suddenly.

Dr. Willy Aspinall (Chief Scientist), Dr.Brian Baptie and Miss Kate Moore left the island today. Professor Stephen Sparks has arrived today to take over as Chief scientist. Mr Colin Walker arrived to continue monitoring of air quality.

Montserrat Volcano Observatory